OpenStudy (anonymous):

If interest is compounded continuosly, after t years the amount in the account is A=Pe^rt (where A =total amount, P=origanal amount invested, and r=interest rate). How long does it take for your investment to double at 4% interest? round to the nearest year9 if your original amount is x, the doubled amount is...

OpenStudy (anonymous):

A = 2P , r = 0.04 , t =? 2P = Pe^rt 2 = e^0.04t ln2=0.04t lne t= ln2/0.04 t = 17.32